A transparent microscope glass (2 x 5 cm) is covered with a TiO2-film. The TiO2-film is placed at the bottom of a Petri dish. The Petri dish is covered with a glass lid transparent to light at wavelengths above 320 nm. When using the lid, oxygen intrusion into the Petri dish can be neglected. The Petri dish is filled with 20 ml pure water (oxygen-free). The Petri dish with the photocatalyst at the bottom is illuminated with continuous light at 360 nm. The intensity of the incident light is 2 mW/cm2. During the illumination the concentration of H2O2 is monitored as a function of time. The concentration reaches a constant level of 2.4 micromolar after 10 minutes. H2O2 is catalytically decomposed on TiO2. The heterogeneous rate constant for the catalytic decomposition is 1.7 x 10-5 m/s. In a separate experiment the light transmission of the TiO2-film at 360 nm was determined to be 5.9 %. Calculate the quantum yield for photocatalytic formation of H2O2 under the present circumstances.


1
Expert's answer
2021-09-21T01:54:15-0400

360-320 = 40nm

20ml

2× 5cm = 10cm

"1.7 x 10^{-5} × 10 × 40×\frac{ 5.9}{100}"

= 0.4012